OpenStudy (anonymous):

solve the equation. 5n-7=28

OpenStudy (khally92):

collect the like terms

OpenStudy (mertsj):

Add 7 to both sides and then divide both sides by 5

OpenStudy (khally92):

like terms are -7 and 28

OpenStudy (anonymous):

Add 7 to 28. 5n=35 -- -- 5 5 n= 7

OpenStudy (anonymous):

solve for n \[5n-7=28\] \[5n=28+7\] \[5n=35\] \[\frac{5n}{5}=\frac{35}{5}\] \[n=7\]